OpenStudy (anonymous):

Perform the indicated problem, and if possible, simplify. show all work. 5/11 + 1/4

OpenStudy (chriss):

first you want to find a common denominator between the two fractions. You can usually start by multiplying each fraction by the denominator/denominator of the other fraction, though often there is a smaller denominator that can be shared between the two fractions.

OpenStudy (anonymous):

Find common denominators. Common denominator is 44. So multiply each fraction by the number that achieves a fraction with denominator of 44. So... (5/11)*(4/4) and (1/4)*(11/11). now you have (20/44)+(11/44)

OpenStudy (chriss):

\[\frac{5}{11}\frac{4}{4}+\frac{1}{4}\frac{11}{11}\]

OpenStudy (anonymous):

I got it that far, I'm at 9/44 I dont think is the lowest it can go...

OpenStudy (chriss):

I don't think your arithmetic is right if you got 9/44. I get something else when I add 20/44+11/44.

OpenStudy (anonymous):

haha i subtracted instead of adding. Thank you!

OpenStudy (chriss):

you're welcome :)